# Visual Autoregressive Modeling:  Scalable Image Generation via Next-Scale Prediction

### v1

@@ Abstract
We present Visual AutoRegressive modeling (VAR), a new generation paradigm that redefines the autoregressive learning on images as coarse-to-fine “next-scale prediction” or “next-resolution prediction”, diverging from the standard raster-scan “next-token prediction”. This simple, intuitive methodology allows autoregressive (AR) transformers to learn visual distributions fast and can generalize well: VAR, for the first time, makes GPT-style AR models surpass diffusion transformers in image generation. On ImageNet 256$\times$256 benchmark, VAR significantly improve AR baseline by improving Fréchet inception distance (FID) from 18.65 to 1.80, inception score (IS) from 80.4 to 356.4, with 20$\times$ faster inference speed. It is also empirically verified that VAR outperforms the Diffusion Transformer (DiT) in multiple dimensions including image quality, inference speed, data efficiency, and scalability. Scaling up VAR models exhibits clear power-law scaling laws similar to those observed in LLMs, with linear correlation coefficients near $-0.998$ as solid evidence. VAR further showcases zero-shot generalization ability in downstream tasks including image in-painting, out-painting, and editing. These results suggest VAR has initially emulated the two important properties of LLMs: Scaling Laws and zero-shot generalization. We have released all models and codes to promote the exploration of AR/VAR models for visual generation and unified learning.
@@ 1 Introduction
The advent of GPT series [48, 49, 9, 45, 1] and other autoregressive (AR) large language models (LLMs) [13, 3, 27, 63, 64, 70, 60, 4, 61] has heralded a new epoch in the field of artificial intelligence. These models exhibit promising intelligence in generality and versatility that, despite issues like hallucinations [28], are still considered to take a solid step toward the general artificial intelligence (AGI). The crux behind these large models is a self-supervised learning strategy – predicting the next token in a sequence, a simple yet profound approach. Studies into the success of these large AR models have highlighted their scalability and generalizabilty: the former, as exemplified by scaling laws [30, 22], allows us to predict large model’s performance from smaller ones and thus guides better resource allocation, while the latter, as evidenced by zero-shot and few-shot learning [49, 9], underscores the unsupervised-trained models’ adaptability to diverse, unseen tasks. These properties reveal AR models’ potential in learning from vast unlabeled data, encapsulating the essence of “AGI”.
@@
Figure 3: Scaling behavior of different models on ImageNet 256$\times$256 conditional generation benchmark. The FID of the validation set serves as a reference lower bound (1.78). VAR with 2B parameters reaches FID 1.80, surpassing L-DiT with 3B or 7B parameters.
Autoregressive modeling requires defining the order of data. Our work reconsiders how to “order” an image. Humans typically perceive or create images in a hierachical manner, first capturing the global structure and then local details. This multi-scale, coarse-to-fine method naturally suggests an “order” for images. Also inspired by the widespread multi-scale designs [40, 39, 62, 31, 33], we define autoregressive learning for images as “next-scale prediction” in Fig. 2 (c), diverging from the conventional “next-token prediction” in Fig. 2 (b). Our approach begins by encoding an image into multi-scale token maps. The autoregressive process is then started from the 1$\times$1 token map, and progressively expands in resolution: at each step, the transformer predicts the next higher-resolution token map conditioned on all previous ones. We refer to this methodology as Visual AutoRegressive (VAR) modeling.
VAR directly leverages GPT-2-like transformer architecture [49] for visual autoregressive learning. On the ImageNet 256$\times$256 benchmark, VAR significantly improves its AR baseline, achieving a Fréchet inception distance (FID) of 1.80 and an inception score (IS) of 356.4, with inference speed 20$\times$ faster (see Sec. 4.4 for details). Notably, VAR surpasses the Diffusion Transformer (DiT) – the foundation of leading diffusion systems like Stable Diffusion 3.0 and SORA [18, 8] – in FID/IS, data efficiency, inference speed, and scalability. VAR models also exhibit scaling laws akin to those witnessed in LLMs. Lastly, we showcase VAR’s zero-shot generalization capabilities in tasks like image in-painting, out-painting, and editing. In summary, our contributions to the community include:
@@ 2.1 Properties of large autoregressive language models
Scaling laws.  Power-law scaling laws [22, 30] mathematically describe the relationship between the growth of model parameters, dataset sizes, computational resources, and the performance improvements of machine learning models, conferring several distinct benefits. First, they facilitate the extrapolation of a larger model’s performance through the scaling up of model size, data size, and computational cost. This helps save unnecessary costs and provides principles to allocate the training budget. Second, the scaling laws have evidenced a consistent and non-saturating increase in performance, corroborating their sustained advantage in enhancing model capability. Propelled by the principles of scaling laws in neural language models [30], several Large Language Models [9, 76, 70, 27, 63, 64] have been proposed, embodying the principle that increasing the scale of models tends to yield enhanced performance outcomes. GPT [49, 9], predicated on a transformer decoder architecture, undergoes generative pre-training and scales the model size to an unprecedented 175B parameters. LLama [63, 64] release a collection of pre-trained and fine-tuned large language models (LLMs) ranging in scale from 7 billion to 70 billion parameters. The manifest efficacy of scaling laws applied to language models has proffered a glimpse into the promising potential of upscale in visual models [5].
Zero-shot generalization.  Zero-shot generalization [55] refers to the ability of a model, particularly a Large Language Model, to perform tasks that it has not been explicitly trained on. Within the realm of the vision community, there is a burgeoning interest in the zero-shot and in-context learning abilities of foundation models, CLIP [47], SAM [35], Dinov2 [44]. Innovations like Painter [69] and LVM [5] have leveraged visual prompts to design in-context learning paradigms, thereby facilitating the generalization to downstream unseen tasks.
@@ 2.2 Visual generation
Image tokenizer and autoregressive models.  Language models [15, 49, 63] rely on Byte Pair Encoding (BPE [20]) or WordPiece algorithms for text tokenization. Visual generation models based on language models also necessitate the encoding of 2D images into 1D token sequences. Early endeavors VQVAE [65] have demonstrated the ability to represent images as discrete tokens, although the reconstruction quality was relatively moderate. VQGAN [19] advances VQVAE by incorporating adversarial loss and perceptual loss to improve image fidelity, and employs a decoder-only transformer to generate image tokens in standard raster-scan autoregressive manner. VQVAE-2 [52] and RQ-Transformer [37] also follow VQGAN’s raster-scan autoregressive method but further improve VQVAE via extra scales or stacked codes. Parti [72] capitalizes on the foundational architecture of ViT-VQGAN [71] to scale the transformer model size to 20 billion parameters, achieving remarkable results in text-to-image synthesis.
Masked-prediction model.  MaskGIT [11] employs a masked prediction framework [15, 6, 21] alongside a VQ autoencoder to generate image tokens based through a “greedy” algorithm. MagViT [73] adapts this approach to video data, and MagViT-2 [74] enhances MaskGIT by introducing an improved VQVAE. MUSE [10] scales MaskGIT’s architecture to 3 billion parameters and merges it with the T5 language model [50], setting new benchmarks in text-to-image synthesis.
Diffusion models  [23, 59, 16, 53] are considered the forefront of visual synthesis, given their superior generation quality and diversity. Progress in diffusion models has centered around improved sampling techniques[26], faster sampling [58, 41], and architectural enhancements [53, 24, 54, 46]. Imagen [54] incorporates the T5 language model [50] for text condition and builds an image generation system through multiple independent diffusion models for cascaded generation and super resolution. Latent Diffusion Models (LDM)[53] apply diffusion in latent space, improving efficiency in training and inference. DiT[46] replaces the traditional U-Net with a transformer-based architecture [66, 17], and is used in recent image or video synthesis systems like Stable Diffusion 3.0 [18] and SORA [8].
@@ 3.1 Preliminary: autoregressive modeling via next-token prediction
Formulation.  Consider a sequence of discrete tokens $x=(x_{1},x_{2},\dots,x_{T})$, where each token $x_{t}\in[V]$ is an integer from a vocabulary of size $V$. The next-token autoregressive model posits that the probability of observing the current token $x_{t}$ depends only on its prefix $(x_{1},x_{2},\dots,x_{t-1})$. This assumption of unidirectional token dependency allows us to decompose the likelihood of sequence $x$ into the product of $T$ conditional probabilities:
@@
Training an autoregressive model $p_{\theta}$ parameterized by $\theta$ involves optimizing the $p_{\theta}(x_{t}\mid x_{1},x_{2},\dots,x_{t-1})$ across a dataset. This optimization process is known as the “next-token prediction”, and the trained $p_{\theta}$ can generate new sequences.
@@
Figure 4: VAR involves two separated training stages. Left: a multi-scale quantized autoencoder (VQVAE) encode an image into $K$ token maps $R=(r_{1},r_{2},\dots,r_{K})$ and is trained by a compound reconstruction loss (5). “Embedding” in the figure means converting discrete tokens into continuous embedding vectors. Right: a VAR transformer is trained by maximizing the likelihood or minimizing the cross-entropy loss in (6) via next-scale prediction. the trained VQVAE is used to produce token map ground truth $R=(r_{1},r_{2},\dots,r_{K})$ for VAR model.
Discussion.  The above tokenizing and flattening enable next-token autoregressive learning on images, but they introduce several issues:
@@
    Mathematical premise violation.  Image encoders typically produce image feature map $f$ with inter-dependent feature vectors $f^{(i,j)}$ for all $i,j$. So after quantization and flattening, the sequence of tokens $(x_{1},x_{2},\dots,x_{h\times w})$ exhibits bidirectional correlations. This contradicts the unidirectional dependency assumption of autoregressive models, which dictates that each token $x_{t}$ should only depend on its prefix $(x_{1},x_{2},\dots,x_{t-1})$.
@@
    Structural degradation.  The flattening disrupts the spatial locality inherent in image feature maps. For instance, the token $q^{(i,j)}$ and its 4 immediate neighbors $q^{(i\pm 1,j)}$, $q^{(i,j\pm 1)}$ are closely correlated due to their proximity. This spatial relationship is compromised in the linear sequence $x$, where unidirectional constraints diminish these correlations.
@@
The disruption of spatial locality (issue 2) is obvious. Regarding issue 1, we present empirical evidence in the Appendix, analyzing the token dependencies in the popular quantized autoencoder [19] and revealing significant bidirectional correlations. The proof of computational complexity for issue 3 is detailed in the Appendix. These theoretical and practical limitations call for a rethinking of autoregressive models in the context of image generation.
@@ 3.2 Visual autoregressive modeling via next-scale prediction
where each autoregressive unit $r_{k}\in[V]^{h_{k}\times w_{k}}$ is the token map at scale $k$, and the sequence $(r_{1},r_{2},\dots,r_{k-1})$ serves as the the “prefix” for $r_{k}$. During the $k$-th autoregressive step, all distributions over the $h_{k}\times w_{k}$ tokens in $r_{k}$ are inter-dependent and will be generated in parallel, conditioned on $r_{k}$’s prefix and associated $k$-th position embedding map. This “next-scale prediction” methodology is what we define as visual autoregressive modeling (VAR), depicted on the right side of Fig. 4.
@@
    The mathematical premise is satisfied if we constrain each $r_{k}$ to depend only on its prefix, that is, the process of getting $r_{k}$ is solely related to $(r_{1},r_{2},\dots,r_{k-1})$. This constraint is acceptable as it aligns with the natural, coarse-to-fine progression characteristics like human visual perception and artistic drawing. Further details are provided in the Tokenization below.
@@ 11 Return: reconstructed image $\hat{im}$;
3.3 Implementation details
VAR tokenizer.  As aforementioned, we use the vanilla VQVAE architecture [19] with a multi-scale quantization scheme with $K$ extra convolutions (0.03M extra parameters). We use a shared codebook for all scales with $V=4096$ and a latent dim of 32. Following the baseline [19], our tokenizer is also trained on OpenImages [36] with the compound loss (5). See the Appendix for more details.
VAR transformer.  Our main focus is on VAR algorithm so we keep a simple model architecture design. We adopt the architecture of standard decoder-only transformers akin to GPT-2 and VQGAN [49, 19], with the only modification of substituting traditional layer normalization for adaptive normalization (AdaLN) – a choice motivated by its widespread adoption and proven effectiveness in visual generative models [33, 34, 32, 57, 56, 29, 46, 12]. For class-conditional synthesis, we use the class embedding as the start token [s] and also the condition of AdaLN. We do not use advanced techniques in modern large language models, such as rotary position embedding (RoPE), SwiGLU MLP, or RMS Norm [63, 64]. Our model shape hyperparameter follows a simple rule [30] that the width $w$, head counts $h$, and drop rate $dr$ are linearly scaled with the depth $d$ as follows:
@@ 3.3 Implementation details
All models are trained with the similar settings: a base learning rate of $10^{-4}$ per 256 batch size, an AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\text{decay}=0.05$, a batch size from 512 to 1024 and training epochs from 200 to 350 (depends on model size). Our subsequent evaluations in Sec. 4 will suggest that such a simple model design are capable of scaling and generalizing well.

### v2

@@ Abstract
We present Visual AutoRegressive modeling (VAR), a new generation paradigm that redefines the autoregressive learning on images as coarse-to-fine “next-scale prediction” or “next-resolution prediction”, diverging from the standard raster-scan “next-token prediction”. This simple, intuitive methodology allows autoregressive (AR) transformers to learn visual distributions fast and can generalize well: VAR, for the first time, makes GPT-style AR models surpass diffusion transformers in image generation. On ImageNet 256$\times$256 benchmark, VAR significantly improve AR baseline by improving Fréchet inception distance (FID) from 18.65 to 1.73, inception score (IS) from 80.4 to 350.2, with 20$\times$ faster inference speed. It is also empirically verified that VAR outperforms the Diffusion Transformer (DiT) in multiple dimensions including image quality, inference speed, data efficiency, and scalability. Scaling up VAR models exhibits clear power-law scaling laws similar to those observed in LLMs, with linear correlation coefficients near $-0.998$ as solid evidence. VAR further showcases zero-shot generalization ability in downstream tasks including image in-painting, out-painting, and editing. These results suggest VAR has initially emulated the two important properties of LLMs: Scaling Laws and zero-shot generalization. We have released all models and codes to promote the exploration of AR/VAR models for visual generation and unified learning.
@@ 1 Introduction
The advent of GPT series [65, 66, 15, 62, 1] and more autoregressive (AR) large language models (LLMs) [22, 4, 38, 82, 83, 90, 78, 5, 79] has heralded a new epoch in the field of artificial intelligence. These models exhibit promising intelligence in generality and versatility that, despite issues like hallucinations [39], are still considered to take a solid step toward the general artificial intelligence (AGI). At the core of these models is a self-supervised learning strategy – predicting the next token in a sequence, a simple yet profound approach. Studies into the success of these large AR models have highlighted their scalability and generalizabilty: the former, as exemplified by scaling laws [43, 35], allows us to predict large model’s performance from smaller ones and thus guides better resource allocation, while the latter, as evidenced by zero-shot and few-shot learning [66, 15], underscores the unsupervised-trained models’ adaptability to diverse, unseen tasks. These properties reveal AR models’ potential in learning from vast unlabeled data, encapsulating the essence of “AGI”.
@@
Figure 3: Scaling behavior of different model families on ImageNet 256$\times$256 generation benchmark. The FID of the validation set serves as a reference lower bound (1.78). VAR with 2B parameters reaches an FID of 1.73, surpassing L-DiT with 3B or 7B parameters.
Autoregressive modeling requires defining the order of data. Our work reconsiders how to “order” an image: Humans typically perceive or create images in a hierachical manner, first capturing the global structure and then local details. This multi-scale, coarse-to-fine nature suggests an “order” for images. Also inspired by the widespread multi-scale designs [54, 52, 81, 44], we define autoregressive learning for images as “next-scale prediction” in Fig. 2 (c), diverging from the conventional “next-token prediction” in Fig. 2 (b). Our approach begins by encoding an image into multi-scale token maps. The autoregressive process is then started from the 1$\times$1 token map, and progressively expands in resolution: at each step, the transformer predicts the next higher-resolution token map conditioned on all previous ones. We refer to this methodology as Visual AutoRegressive (VAR) modeling.
VAR directly leverages GPT-2-like transformer architecture [66] for visual autoregressive learning. On the ImageNet 256$\times$256 benchmark, VAR significantly improves its AR baseline, achieving a Fréchet inception distance (FID) of 1.73 and an inception score (IS) of 350.2, with inference speed 20$\times$ faster (see Sec. 7 for details). Notably, VAR surpasses the Diffusion Transformer (DiT) – the foundation of leading diffusion systems like Stable Diffusion 3.0 and SORA [29, 14] – in FID/IS, data efficiency, inference speed, and scalability. VAR models also exhibit scaling laws akin to those witnessed in LLMs. Lastly, we showcase VAR’s zero-shot generalization capabilities in tasks like image in-painting, out-painting, and editing. In summary, our contributions to the community include:
@@ 2.1 Properties of large autoregressive language models
Scaling laws  are found and studied in autoregressive language models [43, 35], which describe a power-law relationship between the scale of model (or dataset, computation, etc.) and the cross-entropy loss value on the test set. Scaling laws allow us to directly predict the performance of a larger model from smaller ones [1], thus guiding better resource allocation. More pleasingly, they show that the performance of LLMs can scale well with the growth of model, data, and computation and never saturate, which is considered a key factor in the success of  [15, 82, 83, 98, 90, 38]. The success brought by scaling laws has inspired the vision community to explore more similar methods for multimodality understanding and generation [53, 2, 88, 27, 96, 77, 21, 23, 41, 31, 32, 80, 87].
Zero-shot generalization.  Zero-shot generalization [72] refers to the ability of a model, particularly a Large Language Model, to perform tasks that it has not been explicitly trained on. Within the realm of the computer vision, there is a burgeoning interest in the zero-shot and in-context learning abilities of foundation models, CLIP [64], SAM [48], Dinov2 [61]. Innovations like Painter [89] and LVM [6] extend visual prompters [40, 11] to achieve in-context learning in vision.
@@ 2.2 Visual generation
Raster-scan autoregressive models  for visual generation necessitate the encoding of 2D images into 1D token sequences. Early endeavors [20, 84] have shown the ability to generate RGB (or grouped) pixels in the standard row-by-row, raster-scan manner. [69] extends [84] by using multiple independent trainable networks to do super-resolution repeatedly. VQGAN [30] advances [20, 84] by doing autoregressive learning in the latent space of VQVAE [85]. It employs GPT-2 decoder-only transformer to generate tokens in the raster-scan order, like how ViT [28] serializes 2D images into 1D patches. VQVAE-2 [68] and RQ-Transformer [50] also follow this raster-scan manner but use extra scales or stacked codes. Parti [93], based on the architecture of ViT-VQGAN [92], scales the transformer to 20B parameters and works well in text-to-image synthesis.
Masked-prediction model.  MaskGIT [17] employs a VQ autoencoder and a masked prediction transformer similar to BERT [25, 10, 34] to generate VQ tokens through a greedy algorithm. MagViT [94] adapts this approach to videos, and MagViT-2 [95] enhances [17, 94] by introducing an improved VQVAE for both images and videos. MUSE [16] further scales MaskGIT to 3B parameters.
Diffusion models ’ progress has centered around improved learning or sampling [76, 75, 55, 56, 7], guidance [37, 60], latent learning [70], and architectures [36, 63, 71, 91]. DiT and U-ViT [63, 8] replaces or integrates the U-Net with transformer, and inspires recent image [19, 18] or video synthesis systems [12, 33] including Stable Diffusion 3.0 [29], SORA [14], and Vidu [9].
@@ 3.1 Preliminary: autoregressive modeling via next-token prediction
Formulation.  Consider a sequence of discrete tokens $x=(x_{1},x_{2},\dots,x_{T})$, where $x_{t}\in[V]$ is an integer from a vocabulary of size $V$. The next-token autoregressive posits the probability of observing the current token $x_{t}$ depends only on its prefix $(x_{1},x_{2},\dots,x_{t-1})$. This unidirectional token dependency assumption allows for the factorization of the sequence $x$’s likelihood:
@@
Training an autoregressive model $p_{\theta}$ involves optimizing $p_{\theta}(x_{t}\mid x_{1},x_{2},\dots,x_{t-1})$ over a dataset. This is known as the “next-token prediction”, and the trained $p_{\theta}$ can generate new sequences.
@@
Discussion on the weakness of vanilla autoregressive models.  The above approach of tokenizing and flattening enable next-token autoregressive learning on images, but introduces several issues:
@@
    Mathematical premise violation.  In quantized autoencoders (VQVAEs), the encoder typically produces an image feature map $f$ with inter-dependent feature vectors $f^{(i,j)}$ for all $i,j$. So after quantization and flattening, the token sequence $(x_{1},x_{2},\dots,x_{h\times w})$ retains bidirectional correlations. This contradicts the unidirectional dependency assumption of autoregressive models, which dictates that each token $x_{t}$ should only depend on its prefix $(x_{1},x_{2},\dots,x_{t-1})$.
@@
    Inability to perform some zero-shot generalization.  Similar to issue 1), The unidirectional nature of image autoregressive modeling restricts their generalizability in tasks requiring bidirectional reasoning. E.g., it cannot predict the top part of an image given the bottom part.
@@
    Structural degradation.  The flattening disrupts the spatial locality inherent in image feature maps. For example, the token $q^{(i,j)}$ and its 4 immediate neighbors $q^{(i\pm 1,j)}$, $q^{(i,j\pm 1)}$ are closely correlated due to their proximity. This spatial relationship is compromised in the linear sequence $x$, where unidirectional constraints diminish these correlations.
  4. 4)
@@
Issues 2) and 3) are evident (see examples above). Regarding issue 1), we present empirical evidence in Appendix A. The proof of issue 3) is detailed in Appendix B. These theoretical and practical limitations call for a rethinking of autoregressive models in the context of image generation.
[…]
@@ 3.2 Visual autoregressive modeling via next-scale prediction
where each autoregressive unit $r_{k}\in[V]^{h_{k}\times w_{k}}$ is the token map at scale $k$ containing $h_{k}\times w_{k}$ tokens, and the sequence $(r_{1},r_{2},\dots,r_{k-1})$ serves as the the “prefix” for $r_{k}$. During the $k$-th autoregressive step, all distributions over the $h_{k}\times w_{k}$ tokens in $r_{k}$ will be generated in parallel, conditioned on $r_{k}$’s prefix and associated $k$-th position embedding map. This “next-scale prediction” methodology is what we define as visual autoregressive modeling (VAR), depicted on the right side of Fig. 4. Note that in the training of VAR, a block-wise causal attention mask is used to ensure that each $r_{k}$ can only attend to its prefix $r_{\leq k}$. During inference, kv-caching can be used and no mask is needed.
@@
    The mathematical premise is satisfied if we constrain each $r_{k}$ to depend only on its prefix, that is, the process of getting $r_{k}$ is solely related to $r_{\leq k}$. This constraint is acceptable as it aligns with the natural, coarse-to-fine progression characteristics like human visual perception and artistic drawing (as we discussed in Sec. 1). Further details are provided in the Tokenization below.
@@ 11 Return: reconstructed image $\hat{im}$;
4 Implementation details
VAR tokenizer.  As aforementioned, we use the vanilla VQVAE architecture [30] and a multi-scale quantization scheme with $K$ extra convolutions (0.03M extra parameters). We use a shared codebook for all scales with $V=4096$. Following the baseline [30], our tokenizer is also trained on OpenImages [49] with the compound loss (5) and a spatial downsample ratio of 16×16$\times$16 ×.
VAR transformer.  Our main focus is on VAR algorithm so we keep a simple model architecture design. We adopt the architecture of standard decoder-only transformers akin to GPT-2 and VQGAN [66, 30] with adaptive normalization (AdaLN), which has widespread adoption and proven effectiveness in many visual generative models [46, 47, 45, 74, 73, 42, 63, 19]. For class-conditional synthesis, we use the class embedding as the start token [s] and also the condition of AdaLN. We found normalizing $queries$ and $keys$ to unit vectors before attention can stablize the training. We do not use advanced techniques in large language models, such as rotary position embedding (RoPE), SwiGLU MLP, or RMS Norm [82, 83]. Our model shape follows a simple rule like [43] that the width $w$, head counts $h$, and drop rate $dr$ are linearly scaled with the depth $d$ as follows:
@@ 4 Implementation details
All models are trained with the similar settings: a base learning rate of $10^{-4}$ per 256 batch size, an AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\text{decay}=0.05$, a batch size from 768 to 1024 and training epochs from 200 to 350 (depends on model size). The evaluations in Sec. 5 suggest that such a simple model design are capable of scaling and generalizing well.
5 Empirical Results
This section first compares VAR with other image generative model families in Sec. 5.1. Evaluations on the scalability and generalizability of VAR models are presented in Sec. 5.2 and Appendix 6. For implementation details and ablation study, please see Appendix 4 and Appendix 7.